3n - m/2 + 5 = 8n
Solve for n.

Answer by TimothyLamb(4379)

8n - 3n = -m/2 + 5
5n = -m/2 + 5
n = (-m/2 + 5)/5
n = -m/10 + 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php